what is n p r what is n factorial ? what is n c r plz explain their usage and meaning along with their formulae? give some examples of questions with their usage

what is  npr                             what is  n factorial ?


 




what is ncr     plz explain their usage and meaning along with their formulae?


                                                    give some examples  of questions with their usage


 

Grade:12th pass

2 Answers

Sudheesh Singanamalla
13 years ago

nPr is the number of Permutations of 'n' items taken 'r' at a time.

nCr is the number of Combinations of 'n' items taken 'r' at a time.

nPr = n! / (n-r)!

nCr = n! / [ (n-r)! * r! ]

n factorial is the product of n terms and the number of terms preceding it

suppose 'n' is  5 then 5! = 5*4*3*2*1 = 120

similarly n! = n*(n-1)*(n-2)*(n-3)*.......................*3*2*1

nCr is the number of combinations possible .

eg. there are 52 cards in a well shuffled deck. how many combinations are there in which a player can draw only 2 cards at a time ?

nPr is the number of permutations (possbilities).

 

HERE ARE SOME QUESTIONS YOU CAN PRACTICE.

1. There are four bus routes between A and B; and three bus routes
between B and C. A man can travel round-trip in number of ways by bus from A to C
via B. If he does not want to use a bus route more than once, in how many ways can
he make round trip?

2. How many numbers are there between 99 and 1000 having 7 in the units place?

 

Good Luck in your IIT-JEE preparation.

Please approve the answer.

 

vikas askiitian expert
13 years ago

dear nikhil

nPr means out of n population choose r elements and permute them. That is, choose r elements from n and give the total numbers of possibles orders there are.

nCr means out of n population choose r elements. That is, choose r elements from n and do not consider the different orders possible for the same choices of numbers.

So; nPr = nCr*r!